OpenStudy (anonymous):

How many ways are there to choose 8 coins from a piggy bank containing 100 identical pennies and 80 identical nickels. Anybody help me, please

OpenStudy (amistre64):

i recall a formula; something like: n!/(r1! r2! r3! ...)

OpenStudy (amistre64):

but thats how many ways you can pick 8 coins ....

OpenStudy (anonymous):

I have to use generating function to solve the problem. but my answer is incorrect compared to the book.. how can i get 9 with a bunch of factorial

OpenStudy (anonymous):

using "star and bar" method leads to worse result. I stuck.

OpenStudy (amistre64):

pppppppp pppppppn ppppppnp ..... this grouping pattern follows after a row in pascals triangle 1 8 28 56 70 56 28 8 1 but each group of "k" is counted the same; this leads me to: 1 1 1 1 1 1 1 1 1 distinct groups; or 9 groups

OpenStudy (amistre64):

recall that the sets: {a,b,c} and {b,c,a} are different permutations, but are the same combinations

OpenStudy (anonymous):

so?

OpenStudy (amistre64):

youll have to give me something more to go on than "so?" ....

OpenStudy (anonymous):

sorry, I cannot see any link between them. in (a +b)^8 you have 9 terms. but the way we count is combination between 180 identical "things"

OpenStudy (anonymous):

my question is why they have to separate pennies and nickels when both them can be counted as "coin"

OpenStudy (amistre64):

you are being as how many different ways there are to create a group of 8 knowing that we have alot of pennies and nickels. im not sre how the math works it but: there is 1 way to collect 8 pennies and 0 nickels there is 1 way to collect 7 pennies and 1 nickels there is 1 way to collect 6 pennies and 2 nickels there is 1 way to collect 5 pennies and 3 nickels .... etc

OpenStudy (anonymous):

oh oh, i guess, you have (nickel +pennies)^8 and get how many combination between them

OpenStudy (amistre64):

there is alot of extra distraction information in the problem

OpenStudy (amistre64):

with the given information we could delve into the probability of choosing a particular group and so forth; but thats just not relevant information needed for the solution :)

OpenStudy (anonymous):

I have idea and need you help to develop it. are you willing?

OpenStudy (amistre64):

wrong time of day :) ihave 3 class starting in 10 minutes that i need to get to. but id be happy to review your idea when i get back ... which is in about 5 hours :/

OpenStudy (anonymous):

ok, I will write out all and you review later

OpenStudy (amistre64):

have fun, good luck ;)

OpenStudy (anonymous):

take a look at my attached file

OpenStudy (ja1):

isn't it nine?

OpenStudy (amistre64):

i think that you are reading to much into the question; its not considering that each penny (or nickel) is a distinct element that must be counted as a separate element. It says that all the pennies are identical, and all the nickels are identical; and that there are enough elements of each type such that randomly picking out 8 coins would result in how many different groups of coins could be made? It is not asking for probability of picking any one group over the other, its not asking you to classify one penny as different from another, or any other thing. How many ways is there to grab 8 coins from the collection? 8p and 0n 7p and 1n 6p and 2n 5p and 3n 4p and 4n 3p and 5n 2p and 6n 1p and 7n 0p and 8n there are therefore 9 ways in which we can obtain a set of 8 coins of varying combinations of pennies and nickels.

OpenStudy (anonymous):

I got generating function from this problem is \[\frac{ 1 }{(1-x)^2} = \sum_{k=0}^{\infty} (\left(\begin{matrix}k+1 \\ k\end{matrix}\right)x^k\] when x1 + x2 =8 that means I have to figure out the coefficient of x^8. at that moment my k =8 replace to the Combination in the front of x^k I have 9C8 = 9C1 =9. That's is perfect result I get and it is similar to the book. my wonder is the "identical " thing is each of pennies and nickels is identical. that means penny 1 not= penny2 not=penny3 and so on. How can we stop there?

OpenStudy (amistre64):

thats the exact opposite of what it means to be identical ..... if there are 3 bottles of water sitting on the table, same brand, same volume, same same identically same .... and i tell you that i need a bottle of water; does it matter which one you actually pick up? The problem is simplifying the event such that penny1 is identical in properties to penny 2, or to penny k; such that the specific penny you pick up is immaterial. I do not know anything about generating functions at the moment so I cant really comment if thats correct :)

OpenStudy (amistre64):

lets say a chicken is randomly pooping out white eggs and gold eggs; how many different ways are there for a set of 8 eggs to come out?

OpenStudy (amistre64):

thats what im reading at least; im assuming that we arent pulling coins out and then trying to determine if we are replacing them and such ....